OpenStudy (anonymous):

Maria and Edwin have $23.00 each to spend at a book fair, where all students receive a 25% discount. They both want to purchase a copy of the same book, which normally sells for $26.50 plus 10% sales tax. To check if she has enough to purchase the book, Maria takes 25% of $26.50 and subtracts that amount from the normal price. She takes 10% of the discounted selling price and adds it back to find the purchase amount. Edwin takes 75% of the normal purchase price and then computes 110% of the reduced price. Is Maria correct? Is Edwin correct? Do they have enough money to purchase the book? Expla

OpenStudy (anonymous):

hey sorry: i think they are both correct, it is actually the same procedure. First take 75% of 26.50 (same as subtracting 25%) what do you get

OpenStudy (anonymous):

@Maxipoo124 please participate

OpenStudy (anonymous):

ok 75% of 26.50 gives you 19.875 now take 110% 21.86$$$ So they have enough money

OpenStudy (mathstudent55):

They both did the same thing in different forms. Maria: 25% of $26.50 = 0.25 * 26.5 = 6.625 26.5 - 6.625 = 19.875 10% * 19.875 = 1.9875 19.875 + 1.9875 = 21.8625 Rounds off to $21.86 Edwin: 75% of $26.50 = 0.75 * 26.5 = 19.875 Notice that Edwin found the discounter price with a single multiplication. 110% of 19.875 = 1.1 * 19.875 = 21.8625 Rounds off to $21.86 Here, Edwin use a multiplication to add the 10% tax The final prices are the same.